11)Solve the PDE (x+y²)p+yq=z+x².


1
Expert's answer
2021-09-21T02:06:55-0400


this is hyperbolic paraboloid "(x+y\u00b2)p+yq=z+x\u00b2"


"x^2-(x+y\u00b2)p+yq=z"


real solution is;


"0=x^2-px+py^2+yq"